# Attention-Augmented Inverse Reinforcement Learning with Graph Convolutions for Multi-Agent Task Allocation

### v2

@@ Abstract
Multi-agent task allocation (MATA) plays a vital role in cooperative multi-agent systems, with significant implications for applications such as logistics, search and rescue, and robotic coordination. Although traditional deep reinforcement learning (DRL) methods have been shown to be promising, their effectiveness is hindered by a reliance on manually designed reward functions and inefficiencies in dynamic environments. In this paper, an inverse reinforcement learning (IRL)-based framework is proposed, in which multi-head self-attention (MHSA) and graph attention mechanisms are incorporated to enhance reward function learning and task execution efficiency. Expert demonstrations are utilized to infer optimal reward densities, allowing dependence on handcrafted designs to be reduced and adaptability to be improved. Extensive experiments validate the superiority of the proposed method over widely used multi-agent reinforcement learning (MARL) algorithms in terms of both cumulative rewards and task execution efficiency.
@@ I Introduction
With advancements in automatic control, computer technology and artificial intelligence, Multi-Agent Systems (MAS) have attracted increasing attention[1]. As an extensively studied and applied subfield, MAS enables multiple agents to complete tasks cooperatively[2]. In domains such as logistics, search and rescue, the workload and complexity often exceed the capabilities of single agents[3]. MAS address these challenges by offering greater flexibility, fault tolerance, and robustness[4]. However, realizing efficient collaboration in MAS relies heavily on effective task allocation.
@@
Multi-agent task allocation (MATA) is a fundamental issue in MAS research. The quality of task allocation directly affects the system’s operational efficiency. When task complexity exceeds the capability of a single agent, enabling cooperation among agents becomes essential[5]. Practical applications impose diverse demands on task allocation methods. Traditional optimization-based methods such as G-CBBA [6] and ImprovedQPSO [7] have achieved notable success in small-scale, static environments. However, MATA is an NP-hard problem[8] and these methods struggle with scalability, adaptability, and real-time requirements in large-scale, dynamic settings[9].
@@
To address these limitations, deep reinforcement learning (DRL) methods have been applied to MATA. DRL enables agents to learn strategies through interaction with the environment, offering greater adaptability to uncertainty. Unlike traditional methods, DRL exhibits superior adaptability in dynamic and uncertain scenarios[10]. DRL-based methods effectively address the limitations of traditional algorithms, significantly improving task allocation efficiency and robustness.
@@
Despite the progress of DRL in MATA, DRL methods depend heavily on manually designed reward functions, which are often difficult and subjective to define accurately in complex task allocation scenarios. Moreover, agents struggle to balance exploration and exploitation without sufficient prior knowledge[11]. To address these issues, inverse reinforcement learning (IRL) has been explored, particularly in robot path planning. IRL infers reward functions from expert demonstrations, eliminating the need for manual design. Techniques such as Maximum Entropy IRL and Feature Expectation Matching have successfully derived implicit objectives for complex planning tasks[12, 13], enabling robust decision-making in dynamic environments.
@@
Although MATA is more complex—featuring dynamic tasks and inter-agent coordination—it shares core challenges with path planning, such as sequential decision-making and global optimization. Nevertheless, IRL research in MATA remains scarce. Most existing studies focus on algorithmic performance, while the integration of global information and prior knowledge into multi-agent learning remains underexplored[14, 15]. Motivated by the success of IRL in path planning, an IRL-based framework is extended to the MATA setting to address the challenges of reward function design and agent coordination in complex environments. To improve reward modeling and decision-making, multi-head self-attention (MHSA) and graph attention networks (GAT) are incorporated. MHSA is utilized to capture intricate dependencies within agent trajectories, while GAT is employed to model global agent-task relationships through graph-structured state representations. Through these mechanisms, more expressive feature representations and robust reward estimations are enabled, enhancing the overall efficiency and adaptability of task allocation.
@@
    Extensive experiments demonstrate the effectiveness of the proposed framework across different scales. The framework is evaluated under varying task scales with a fixed number of agents and under varying agent scales with a fixed number of tasks. Compared with baseline methods, it achieves significant improvements in reward efficiency and overall coordination. Ablation studies further validate the contribution of each module to the overall performance.
The remainder of the paper is organized as follows. Section II reviews related work. Section III describes the MATA system model and problem formulation. Section IV presents the proposed approaches. Section V describes the experiments, results analysis. Finally, Section VI concludes the paper.
@@ II-A DRL based task allocation
In the CTDE paradigm, decentralized decisions are made after centralized training, offering robustness against communication failures and improved scalability. Various approaches have been developed under this framework. CapAM [19] combines capsule networks with attention-based GNNs to capture local and global features of task graphs, maintaining performance as task-agent scales increase. An auction-based RL framework [14] coordinates agents through proposer-responder roles, enabling fast decision-making in dynamic environments. Hierarchical reinforcement learning [20] decomposes tasks into subchains, using QMIX-based value decomposition to ensure effective collaboration.
To address practical constraints in real-world scenarios, more recent methods have incorporated advanced architectures and application-specific adaptations. LIA-MADDPG [21] enhances robot swarm coordination by aggregating local information during distributed execution. In cooperative transport [22], task priorities are integrated with global communication to support coordination under partial observability. MAGNNET [23] embeds GNNs into the CTDE structure to achieve scalable, conflict-free task assignment among autonomous vehicles. In UAV-assisted mobile crowd sensing [24], task allocation and trajectory planning are jointly optimized using a CTDE-based MARL framework with integrated greedy strategy to enhance coverage efficiency.
@@ IV-A DRL based Task Allocation
State Space $S$: In factory and logistics park scenarios, an individual agent’s perception via vision or LiDAR is limited. However, reliable communication enables agents to share global state information, including the positions of agents and tasks, ensuring equal access to a comprehensive state representation.
@@ IV-B IRL based Task Allocation
The framework proposed in this study consists of a generation component and an adversarial component, as illustrated in Fig. 1. The framework incorporates these key modules: the yellow-shaded area represents the MARL algorithm, the peachy-shaded area corresponds to the baseline generator enhanced by the MHSA mechanism, and the green-shaded area represents the generator further enhanced by a hierarchical MHSA combined with a global graph neural network. These three parts make up the generator. The blue-shaded area represents the discriminator.
@@
Then, the Query, Key and Value required for computing the attention mechanism are defined as: $\mathcal{Q}=EB_{ki}W_{\mathcal{Q}}$, $\mathcal{K}=EB_{ki}W_{\mathcal{K}}$, $\mathcal{V}=EB_{ki}W_{\mathcal{V}}$, where $W_{\mathcal{Q}},W_{\mathcal{K}},W_{\mathcal{V}}\in\mathbb{R}^{d\times d_{head}}$ are learnable linear transformation matrices, and $d_{head}$ is the dimension of a single attention head. For an attention head $\text{head}_{n}$, the output is:
@@
Subsequently, the outputs of all $h$ attention heads are concatenated and fused using a linear layer. The result is the feature matrix $H_{ki}=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h})\cdot W_{O}\in\mathbb{R}^{L\times d}$, where $W_{O}\in\mathbb{R}^{(h\cdot d_{\text{head}})\times d}$ is the final linear transformation matrix for MHSA. Finally, a fully connected neural network (FCN) is used to generate the coefficients $\alpha$ and $\beta$ for modifying the reward function: $[\alpha,\beta]^{T}=W_{r}\cdot\frac{1}{L}\sum_{v=1}^{L}H_{ki}^{(v)}+b_{r}$, where $W_{r}\in\mathbb{R}^{2\times d}$ and $b_{r}\in\mathbb{R}^{2}$ are the projection matrix and bias, respectively. And $H_{ki}^{(v)}\in\mathbb{R}^{d}$ denotes the $v$-th row of the matrix $H_{ki}$, corresponding to the feature representation at position $v$. After completing the generation process, the newly modified reward function is obtained:
@@
By minimizing the binary cross-entropy loss, the discriminator learns to accurately classify expert trajectories while distinguishing them from generated trajectories. This feedback serves as a critical learning signal for the generator, guiding it to iteratively refine its output to better mimic expert behavior. Through this adversarial interaction, the training process is progressively driven toward convergence.
@@ IV-C Global Information Feature Extraction
Traditional IRL primarily relies on agent trajectory data, overlooking global state information in cooperative or competitive scenarios. By integrating global states, the generator can simultaneously capture local behaviors and environmental features, extracting feature representations that better align with the true reward structure. This enhances the consistency with expert behaviors and improves the framework’s generalization capability in complex tasks.
@@
2:     Initialize the actor network, the critic network and the entropy optimizer of agent $\lambda_{i}$ with parameters $\eta^{a}$, $\eta^{c}$, $E_{tar}$
@@ V Experiment and Result
In order to verify the effectiveness and efficiency of the proposed approach, experiments are conducted in the open-source multi-agent environment, which consists of agents and tasks inhabiting a two-dimensional world with continuous space and discrete time. A comparison is conducted between the proposed method and two classical MARL algorithms, namely MASAC [37] and MAPPO [38].
@@ V-A Experiment Setup
To provide expert demonstrations for the IRL framework, we generated expert trajectories using a classical optimization-based task allocation algorithm under a fully static environment. Although such optimization methods are not feasible in dynamic environment, they provide structured, high-quality demonstrations suitable for reward inference. These trajectories were then used to guide the IRL-based learning in more general scenarios.
@@ V-B Performance Analysis
Figure 3: Performance of cumulative reward. (a) Normal reward. (b) Sparse reward. (c) Dense reward.
@@
Under the normal reward condition, our method reaches an average reward of 174.07 with only 74.06 time steps, significantly outperforming MAPPO and MASAC, which require more steps and yield lower rewards. In the sparse reward scenario, where learning is generally more difficult due to limited feedback, the advantage of the proposed approach becomes more evident: it achieves nearly 40% higher reward than MAPPO and almost three times that of MASAC, while also consuming fewer time steps. In the dense reward setting, all methods perform better due to frequent feedback, but our method still maintains the lead with the highest reward of 326.95 and the lowest time consumption of 80.15, demonstrating its ability to balance reward maximization and task efficiency even in richly supervised environments.
To further examine the scalability and coordination capability of the proposed method, additional experiments are conducted under a fixed number of 40 tasks while varying the number of agents among 5, 10, and 20. The results are presented in Fig. 5 for cumulative reward and Fig. 6 for time-step consumption.
@@ V-C Ablation Analysis
To assess the contributions of different components in the proposed framework, ablation experiments are conducted by selectively removing key elements and analyzing their impact on performance. Specifically, the effects of removing global observations, agent trajectory information and expert trajectories along with the generative adversarial framework are examined. The results are shown in Fig. 7.
When global observations are removed, agents rely solely on trajectory information, leading to a decline in cumulative rewards. This indicates that global observations provide essential contextual information for multi-agent coordination, although trajectory-based learning still enables some adaptation.
Removing agent trajectory processing further reduces performance, as policies trained only with expert demonstrations and global observations lack direct access to past actions. This limitation hinders adaptive decision-making, reducing policy flexibility and stability.
The most significant degradation occurs when expert trajectories and the GA framework are removed, forcing the model to rely entirely on environmental rewards. Without imitation learning, training becomes less stable, convergence slows, and suboptimal policies are more likely.
Overall, removing expert trajectories and the generative adversarial framework has the most severe impact, while removing global observations or agent trajectories still permits learning but at reduced efficiency. These findings highlight the fundamental role of GAIL in stabilizing training and improving policy quality, while both global observations and agent trajectories contribute to effective decision-making.
@@ V-D Complexity Analysis
In this study, we propose an IRL-based MATA framework, which integrates trajectory representation learning from GAIL with a global state modeling module based on graph attention networks. Specifically, agent trajectories are encoded using a MHSA mechanism to capture temporal dependencies, while global agent-task relations are modeled through the graph attention mechanism to enhance coordination. These features are jointly processed within a generative adversarial structure to infer reward coefficients, enabling more flexible and informative policy learning without relying on manually designed reward functions. Experimental results validate the superiority of the proposed method over well-recognized baselines in terms of both cumulative reward and task efficiency. In addition, ablation studies confirm the effectiveness of expert trajectories, global observations and adversarial learning in improving learning stability and policy quality. Overall, this framework offers a modular and scalable solution for cooperative multi-agent decision-making, effectively bridging expert knowledge and autonomous learning. To further enhance the generality and applicability of the proposed framework, future work will focus on extending the method to heterogeneous agents and tasks, as well as incorporating dynamic graph structures to better capture real-world communication and coordination constraints.
@@
This work was supported by the National Natural Science Foundation of China under Grant No.62133011. The authors would like to thank TÜV SÜD for the kind and generous support. We are also grateful for the efforts from our colleagues in Sino German Center of Intelligent Systems.

### v3

@@ Abstract
Multi-agent task allocation (MATA) plays a vital role in cooperative multi-agent systems, with significant implications for applications such as autonomous vehicle coordination, intelligent vehicular logistics and cooperative robotics in transportation systems. Although traditional deep reinforcement learning (DRL) methods have been shown to be promising, their effectiveness is hindered by a reliance on manually designed reward functions and inefficiencies in dynamic environments. In this paper, an inverse reinforcement learning (IRL)-based framework is proposed, in which multi-head self-attention (MHSA) and graph attention mechanisms are incorporated to enhance reward function learning and task execution efficiency. Expert demonstrations are utilized to infer optimal reward densities, allowing dependence on handcrafted designs to be reduced and adaptability to be improved. The proposed framework demonstrates notable effectiveness in optimizing cumulative rewards and improving coordination efficiency, as validated through extensive experimentation.
@@ I Introduction
With advancements in automatic control, computer technology and artificial intelligence, Multi-Agent Systems (MAS) have attracted increasing attention[1]. As an extensively studied and applied subfield, MAS enables multiple agents to complete tasks cooperatively[2]. In domains such as logistics, search and rescue, the workload and complexity often exceed the capabilities of single agents[3]. MAS address these challenges by offering greater flexibility, fault tolerance, and robustness[4]. A representative application is warehouse automation using Automated Guided Vehicles (AGVs), where coordinated operation among multiple AGVs is crucial for efficient material handling and distribution. However, realizing efficient collaboration in MAS relies heavily on effective task allocation.
@@
As shown in Fig. 1, multi-agent task allocation (MATA) is a fundamental issue in MAS research. The quality of task allocation directly affects the system’s operational efficiency. When task complexity exceeds the capability of a single agent, enabling cooperation among agents becomes essential[5]. Practical applications impose diverse demands on task allocation methods. Traditional optimization-based methods such as G-CBBA [6] and ImprovedQPSO [7] have achieved notable success in small-scale, static environments. However, MATA is an NP-hard problem[8] and these methods struggle with scalability, adaptability, and real-time requirements in large-scale, dynamic settings[9].
@@
To address these limitations, deep reinforcement learning (DRL) methods have been applied to MATA. DRL allows agents to autonomously acquire policies by interacting with dynamic environments, offering greater adaptability to uncertainty. Unlike traditional methods, DRL exhibits superior adaptability in dynamic and uncertain scenarios[10]. DRL-based methods effectively address the limitations of traditional algorithms, significantly improving multi-agent task coordination in complex and dynamic vehicular systems, enhancing operational efficiency and robustness.
Figure 1: Schematic diagram of MATA for AGV formation.
@@
Despite advancements in DRL, most existing approaches still rely heavily on manually designed reward functions, which are often difficult and subjective to define accurately in complex task allocation scenarios. Moreover, agents struggle to balance exploration and exploitation without sufficient prior knowledge[11]. To address these issues, inverse reinforcement learning (IRL) has been explored, particularly in robot path planning. IRL infers reward functions from expert demonstrations, eliminating the need for manual design. Techniques such as Maximum Entropy IRL and Feature Expectation Matching have successfully derived implicit objectives for complex planning tasks[12, 13], enabling robust decision-making in dynamic environments.
@@
Although MATA is more complex—featuring dynamic tasks and inter-agent coordination—it shares core challenges with path planning, such as sequential decision-making and global optimization. Nevertheless, IRL research in MATA remains scarce. Most existing studies focus on algorithmic performance, while the integration of global information and prior knowledge into multi-agent learning remains underexplored[14, 15].
Motivated by the success of IRL in path planning, an IRL-based framework is extended to the MATA setting to address the challenges of reward function design and agent coordination in complex environments. To improve reward modeling and decision-making, multi-head self-attention (MHSA) and graph attention networks (GAT) are incorporated. MHSA is utilized to capture intricate dependencies within agent trajectories, while GAT is employed to model global agent-task relationships through graph-structured state representations. Through these mechanisms, more expressive feature representations and robust reward estimations are enabled, enhancing the overall efficiency and adaptability of task allocation.
@@
    Extensive experiments demonstrate the effectiveness of the proposed framework across different scales. The framework is evaluated under varying task scales with a fixed number of agents and under varying agent scales with a fixed number of tasks. Experimental results show that the proposed method effectively enhances reward efficiency and coordination quality in MATA.
The remainder of the paper is organized as follows. Section II reviews related work. Section III describes the MATA system model and problem formulation. Section IV presents the proposed approaches. Section V describes the experiments and results. Finally, Section VI concludes the paper.
@@ II-A DRL based task allocation
In the CTDE paradigm, decentralized decisions are made after centralized training, offering robustness against communication failures and improved scalability. Various approaches have been developed under this framework. CapAM [19] combines capsule networks with attention-based GNNs to capture local and global features of task graphs, maintaining performance as task-agent scales increase. An auction-based RL framework [14] coordinates agents through proposer-responder roles, enabling fast decision-making in dynamic environments. Hierarchical reinforcement learning [20] decomposes tasks into subchains, using QMIX-based value decomposition to ensure effective collaboration. To address practical constraints in real-world scenarios, more recent methods have incorporated advanced architectures and application-specific adaptations. LIA-MADDPG [21] enhances robot swarm coordination by aggregating local information during distributed execution. In cooperative transport [22], task priorities are integrated with global communication to support coordination under partial observability. MAGNNET [23] embeds GNNs into the CTDE structure to achieve scalable, conflict-free task assignment among autonomous vehicles. In UAV-assisted mobile crowd sensing [24], task allocation and trajectory planning are jointly optimized using a CTDE-based MARL framework with integrated greedy strategy to enhance coverage efficiency.
@@ III System Model and Problem Formulation
To support efficient and conflict-free task allocation in the MATA setting, key assumptions and constraints on agent behavior, task timing, and energy consumption are defined in the system model. Based on these formulations, an optimization problem is constructed to capture the trade-offs between energy usage and task completion efficiency.
@@ IV-A DRL based Task Allocation
State Space $S$: In smart transportation and vehicular logistics scenarios, an individual agent’s perception via vision or LiDAR is limited. However, reliable communication enables agents to share global state information, including the positions of agents and tasks, ensuring equal access to a comprehensive state representation.
@@ IV-B IRL based Task Allocation
The framework proposed in this study consists of a generation component and an adversarial component, as illustrated in Fig. 2. The framework incorporates these key modules: the yellow-shaded area represents the MARL algorithm, the peachy-shaded area corresponds to the baseline generator enhanced by the MHSA mechanism, and the green-shaded area represents the generator further enhanced by a hierarchical MHSA combined with a global graph neural network. These two parts in the dashed box make up the generator. The blue-shaded area represents the discriminator.
@@
Then, the Query, Key and Value required for computing the attention mechanism are defined as: $\mathcal{Q}=EB_{ki}W_{\mathcal{Q}}$, $\mathcal{K}=EB_{ki}W_{\mathcal{K}}$, $\mathcal{V}=EB_{ki}W_{\mathcal{V}}$, where $W_{\mathcal{Q}},W_{\mathcal{K}},W_{\mathcal{V}}\in\mathbb{R}^{d\times d_{head}}$ are learnable linear transformation matrices applied to the input embeddings, and $d_{head}$ is the dimension of a single attention head. For an attention head $\text{head}_{n}$, the output is:
@@
Subsequently, the outputs of all $h$ attention heads are concatenated and fused using a linear layer. The result is the feature matrix $H_{ki}=\text{Concat}(\text{head}_{1},\text{head}_{2},\dots,\text{head}_{h})\cdot W_{O}\in\mathbb{R}^{L\times d}$, where $W_{O}\in\mathbb{R}^{(h\cdot d_{\text{head}})\times d}$ is the final linear transformation matrix for MHSA. Finally, a fully connected neural network (FCN) is used to generate the coefficients $\alpha$ and $\beta$ for modifying the reward function: $[\alpha,\beta]^{T}=W_{r}\cdot\frac{1}{L}\sum_{v=1}^{L}H_{ki}^{(v)}+b_{r}$, where $W_{r}\in\mathbb{R}^{2\times d}$ and $b_{r}\in\mathbb{R}^{2}$ are the projection matrix and bias, respectively. And $H_{ki}^{(v)}\in\mathbb{R}^{d}$ denotes the $v$-th row of the matrix $H_{ki}$, corresponding to the feature representation at position $v$. After completing the generation process, the modified reward function is obtained:
@@
By minimizing the binary cross-entropy loss, the discriminator learns to accurately classify expert trajectories while distinguishing them from generated trajectories. This feedback serves as a critical learning signal for the generator, guiding it to iteratively refine its output to better mimic expert behavior, as shown by the data paths composed of green arrows in Fig. 2. Through this adversarial interaction, the training process is progressively driven toward convergence.
@@ IV-C Global Information Feature Extraction
Traditional IRL primarily relies on agent trajectory data, overlooking global state information that is often crucial in cooperative or competitive multi-agent scenarios. This limitation can lead to suboptimal policy learning, especially when agent interactions are influenced by shared environmental dynamics. By integrating global states, the generator can simultaneously capture local behaviors and contextual environmental features, enabling the extraction of richer and more discriminative feature representations that better align with the true reward structure. As a result, the learned policy demonstrates enhanced consistency with expert behaviors and improved generalization capability when applied to complex and dynamic tasks.
@@
2:     Initialize the agent with parameters $\eta$
@@ V Experiment and Result
In order to verify the effectiveness and efficiency of the proposed approach, experiments are conducted in the open-source multi-agent environment, which consists of agents and tasks inhabiting a two-dimensional world with continuous space and discrete time. A comparison is conducted between the proposed method and two widely used MARL algorithms, namely MASAC [37] and MAPPO [38].
@@ V-A Experiment Setup
To provide expert demonstrations for the IRL framework, expert trajectories were generated using a classical optimization-based task allocation algorithm under a fully static environment. Although such optimization methods are computationally expensive and thus not feasible in dynamic environments due to real-time constraints and inherent unpredictability, structured and high-quality trajectories suitable for reward inference can still be produced. These trajectories were then used to guide the IRL-based learning in more general scenarios.
@@ V-B Performance Analysis
Under the normal reward condition, our method achieves an average reward of 174.07 with only 74.06 timesteps, significantly outperforming MAPPO and MASAC, both of which require more steps and yield lower rewards. In the sparse reward setting, where limited feedback makes learning more challenging, the advantage becomes even more pronounced: our method delivers nearly 40% higher rewards than MAPPO and almost triple that of MASAC, while consuming fewer timesteps. In the dense reward case, although all methods improve due to frequent feedback, our method still leads with the highest reward of 326.95 and the lowest time usage of 80.15, effectively balancing reward maximization and task efficiency.
@@
To evaluate scalability and coordination, further experiments are conducted with 40 fixed tasks and varying agent numbers (5, 10, 20). Cumulative rewards and time-step consumption are shown in Fig. 6 and Fig. 7, respectively.
@@ V-C Ablation Analysis
To assess the contributions of different components in the proposed framework, ablation experiments are conducted by selectively removing key elements and analyzing their impact on performance. Specifically, the effects of removing global observations, agent trajectory information, and expert trajectories along with the generative adversarial framework are systematically examined. The corresponding results are presented in Fig. 8.
When global observations are excluded, agents must rely solely on local trajectory information, resulting in a noticeable decline in cumulative rewards. This outcome suggests that global observations provide critical contextual cues that facilitate coordinated behavior among agents. Nevertheless, the presence of trajectory-based learning still allows for a degree of adaptation, albeit less effectively.
Further performance degradation is observed when agent trajectory processing is removed. In this case, policies trained only with global observations and expert demonstrations lack direct access to the agents’ own past actions. This absence restricts the policy’s ability to make informed and adaptive decisions over time, thereby reducing both flexibility and stability in dynamic environments.
@@
The most significant performance drop occurs when expert trajectories and the generative adversarial framework are entirely removed, forcing the model to rely exclusively on environmental rewards for training. Without the guidance of imitation learning, the training process becomes less stable, convergence is substantially slower, and the resulting policies tend to be suboptimal.
Overall, the removal of expert trajectories and the generative adversarial framework has the most detrimental effect on performance. In contrast, eliminating global observations or agent trajectory inputs still permits a certain level of learning but with diminished efficiency. These results underscore the foundational importance of the GAIL component in stabilizing training and enhancing policy quality, while also highlighting the complementary roles that global observations and agent trajectories play in supporting effective multi-agent decision-making.
@@ V-D Complexity Analysis
This paper introduces a novel IRL-based MATA framework, which integrates trajectory representation learning from GAIL with a global state modeling module based on graph attention networks. Specifically, agent trajectories are encoded using a MHSA mechanism to capture temporal dependencies, while global agent-task relations are modeled through the graph attention mechanism to enhance coordination. These features are jointly processed within a generative adversarial structure to infer reward coefficients, enabling more flexible and informative policy learning without relying on manually designed reward functions.Experimental results validate the superiority of the proposed method over well-recognized baselines in terms of both cumulative reward and task efficiency. Overall, this framework offers a modular and scalable solution for cooperative multi-agent decision-making, effectively bridging domain expertise with autonomous policy learning in multi-agent systems. To further enhance the generality and applicability of the proposed framework, future work will focus on extending the method to heterogeneous agents and tasks, as well as incorporating dynamic graph structures to better capture real-world vehicular communication constraints and coordination challenges in intelligent transportation and connected vehicle systems.
@@
This work was supported in part by a national science foundation and by an international industry partner. The authors also appreciate the contributions from colleagues involved in the collaborative research center on intelligent systems.